The Target Copy Company is contemplating the replacement of its old printing machine with a new model costing $60,000. The old machine, which originally cost $40,000, has 6 years of expected life remaining and a current book value of $30,000 versus a current market value of $24,000. Target's corporate tax rate is 40 percent. If Target sells the old machine at market value, what is the initial after-tax outlay for the new printing machine?

-$22,180

-$30,000

-$33,600

-$36,000
Answer #1

Solution :

Statement showing calculation of Initial after-tax outlay for the new printing machine:

Sl.No.

Particulars

Amount

1.

Sale value of the Asset ( i.e., Market value )

$ 24,000

2.

Less : Book value of the Asset

$ 30,000

3.

Loss on sale of Asset = (1) – (2)

- $ 6,000

4.

Tax Rate

40 %

5.

Tax savings on loss of sale of asset = ( $ 6,000 * 0.40 )

$ 2,400

6.

After Tax salvage value = Sale value of asset + Tax savings on loss of sale of asset = $ 24,000 + $ 2,400 = (1) + (5 )

$ 26,400

7.

Investment in new model of printing machine

- $ 60,000

8.

Initial after-tax outlay for the new printing machine

= (6) + (7) = $ 26,400 - $ 60,000

- $ 33,600

Thus the Initial after-tax outlay for the new printing machine is = - $ 33,600.

The solution is Option 3 = - $ 33,600